# Adaptive Uncertainty Quantification for Generative AI

### v1

@@ 2.2 Robust Regression Trees
(Conformal Forests) Gasparin and Ramdas [12] have recently proposed a method for combining conformal sets via a weighted majority vote procedure which preserves a weakened coverage guarantee. We briefly describe how this can be used to extend Conformal Tree to the random forest setting. One could grow $h$ different trees, for example, by subsampling the calibration data, randomizing splitting dimensions, or both. The prediction set corresponding to each tree can be combined using the weighted majority vote procedure [12], yielding an aggregated interval. When applied to our robust dyadic trees in Algorithm 2, the aggregated interval will have marginal coverage with probability at least $1-2\alpha-2\delta$. Such an aggregated interval may sacrifice conditional coverage in each leaf node, as well as some of the coverage guarantee, in exchange for a smoother and more locally adaptive conformal set. The empirical performance of our naive implementation of Conformal Forests can be found in Table 1.
@@ 3 Theoretical Underpinnings
Let $(X_{1},Y_{1}),\ldots,(X_{n+1},Y_{n+1})$ be i.i.d. samples from a joint probability measure $\mathbb{P}$. Let $\widehat{\mathfrak{X}}(\mathcal{D}_{1:j})$ denote the partition resulting from Algorithm 2 on data $(X_{1},S(X_{1},Y_{1})),\ldots,$ $(X_{j},S(X_{j},Y_{j})))$ for $j\in\{n,n+1\}$. Denote by $m$ the hyperparameter limiting the minimum number of samples in each leaf node. Define
$\delta(n,m):=\frac{2}{m}+\binom{n+1}{m}\left(\frac{m}{n+1}\right)^{m}\left(1-\frac{m}{n+1}\right)^{n+1-m}$ (3.1)
@@
Under the conditions and notations of Lemma 1, consider the prediction set in (2.1) with $\mathfrak{X}=\widehat{\mathfrak{X}}(\mathcal{D}_{1:{n}})=\{\mathscr{X}_{k}\}_{k=1}^{K}$ and denote it by $\widehat{C}^{\widehat{\mathfrak{X}}(\mathcal{D}_{1:{n}})}_{n,m}(X_{n+1})$. Then, we have for any $k\in\{1,\dots,K\},$
$\mathbb{P}\{Y_{n+1}\in\widehat{C}^{\widehat{\mathfrak{X}}(\mathcal{D}_{1:{n}})}_{n,m}(X_{n+1})\mid X_{n+1}\in\mathscr{X}_{k},\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})\}\geq 1-\alpha-\delta(n,m)\,$
where $\delta(n,m)$ is as defined in (3.1).
@@
Let $(X_{1},Y_{1}),\ldots,(X_{n+1},Y_{n+1})$ be i.i.d. samples from a joint probability measure $\mathbb{P}$. Denote by $\widehat{C}_{n,m}$ the interval constructed by Algorithm 1, with a minimum of $m$ samples in each leaf. Then, we have
@@
$\mathbb{P}\{Y_{n+1}\in\widehat{C}_{n,m}(X_{n+1})\}\leq 1-\alpha+\frac{1}{m-2}+\delta(n,m)\,.$ (3.4)
@@
While the bounds in Theorem 3.2 are not entirely optimistic as we upper-bound the probability of miscoverage by 1 when $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})\neq\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$. We found empirically, however, that the Conformal Tree yields the desired coverage on most examples we tried (as summarized in Table 1). One may also restore the originally intended coverage level $1-\alpha$ by using the $1-\alpha+\delta(n,m)$ quantile (if below one) to define the conformal interval within each leaf region.
@@ 4 Simulation and Benchmark Studies
In this section, we consider regression tasks on simulated and real benchmark data. For regression tasks, we use the absolute residuals as the conformity score, and compare the average length of the prediction interval as our primary evaluation metric.
Figure 1: Motivating example: Conformal Tree can adaptively reflect local changes without the necessity for an auxiliary model of quantiles or conformity scores fit on the training dataset. We display the average width, empirical coverage on held-out test data, and the proportion of test points with a smaller prediction set than that of split conformal.
@@ 4.1 Conformal Tree is Adaptive
These two setups represent heteroskedasticity for distinct reasons. Data 1 has the noise standard deviation scaling with $X$, causing the distribution of residuals to vary across the domain. Therefore, any single regression (mean) model cannot avoid increasing errors for the larger covariate due to the increasing variation of the data. Data 2 is difficult to model near $X=0$ due to the $X^{-3}$ term inside the $\sin$, causing rapid change in the mean function that requires a lot of data to model accurately. Therefore, due to the lack of model fit, the error is large for the smaller covariate values. We apply our Conformal Tree method ($m=20$ and maximum leaves as $8$) and compare it with the standard split-conformal prediction. Figure 1 highlights that Conformal Tree can adaptively reflect local changes without the necessity for an auxiliary model of quantiles or conformity scores on the training dataset. As the partitioning method is robust to the addition of a new iid data point, the obtained prediction intervals with $\alpha=0.1$ still have good coverage as 0.908 for Data 1 and 0.910 for Data 2 when averaged over 10 random trials. For the standard conformal prediction, the average coverage was 0.902 and 0.900, respectively. Furthermore, the average percentage of points for which the interval was narrower is 0.614 for Data 1 and 0.5 for Data 2. Figure 1 corresponds to a single random trial among the 10 random trials.
@@ 4.2 Benchmark Datasets
We consider various real data examples that had previously been used in the literature for comparing the efficacy of regression prediction intervals [29, 31, 27]. The number of observations in these datasets vary from 200 to about 50,000, and the number of predictor variables ranges from 1 to 100. We randomly split each dataset into training, calibration, and test sets, with proportions 0.4, 0.4, and 0.2 respectively. For each dataset, we repeat the entire procedure 10 times, including splitting the data and forming the intervals. We report the average metrics taken over these trials. For the “black-box” predictive model $\hat{\mu}$, we consider a random forest model. Additional information about the datasets, predictive models, and experiments can be found in Supplement C, as well as additional results.
Dataset Split Conformal Conformal Tree Conformal Tree (CART) Conformal Forest Locally Weighted (Forest) CQR
Width Cov. Width Cov. P.B. Width Cov. P.B. Width Cov. P.B. Width Cov. P.B. Width Cov. P.B.
Data 1 4.43 0.91 4.03 0.9 0.66 3.91 0.9 0.62 4.1 0.91 0.58 4.79 0.92 0.56 4.25 0.9 0.66
Data 2 2.83 0.91 3.75 0.92 0.72 3.5 0.91 0.74 3.66 0.91 0.6 4.69 0.91 0.47 3.09 0.9 0.58
bike 1.31 0.9 1.38 0.9 0.52 1.19 0.9 0.59 1.12 0.9 0.73 0.99 0.89 0.74 0.64 0.89 0.91
bio 1.66 0.9 1.64 0.9 0.59 1.6 0.9 0.51 1.67 0.9 0.52 1.28 0.9 0.71 1.16 0.9 0.68
community 1.91 0.9 1.97 0.91 0.46 1.68 0.89 0.68 2.09 0.91 0.61 1.57 0.9 0.68 1.54 0.9 0.68
concrete 0.63 0.89 0.68 0.9 0.57 0.65 0.9 0.59 0.7 0.89 0.48 0.63 0.9 0.59 0.57 0.9 0.65
homes 0.61 0.9 0.64 0.9 0.63 0.62 0.9 0.35 0.7 0.9 0.62 0.58 0.9 0.73 0.56 0.9 0.75
star 0.17 0.9 0.17 0.91 0.59 0.17 0.9 0.44 0.2 0.91 0.37 0.18 0.9 0.45 0.18 0.9 0.4
Table 1: Empirical results on real and synthetic data. We display the average width, coverage, and proportion of test intervals smaller than that of split conformal (P.B is proportion better). The confidence level is set to $\alpha=0.1$ and table cells are means over ten trials. The predictive model for $\hat{\mu}$ is a random forest. Locally-weighted conformal methods use an auxiliary model fit on the training data. CQR uses a quantile regression model fit on the training data.
Table 1 compares the resulting average interval widths arising from conformal prediction methods, where a random forest model is used as our predictive model $\hat{\mu}$. We compare with standard split-conformal prediction [16], locally-weighted conformal inference (LWCI) [16], and conformalized quantile regression (CQR) [27]. Locally-weighted conformal inference is similar to our method in that it relies on a locally varying model for the conformity score $\hat{\sigma}:\mathcal{X}\to\mathbb{R}$, which is fit on the training data. We use a random forest model for $\hat{\sigma}$ in our experiments. We also display a variant of our method that uses standard CART for partitioning, as well as a random forest variant (Remark 2). These methods lack the same theoretical coverage guarantee but cover at the correct level empirically. Table 1 also shows the average empirical coverage on the holdout test data for each method. We observe that each method achieves the desired coverage level of 0.9. We also display the proportion of test points in which the conformal intervals are smaller than that of split conformal.
We observe competitive performance between Conformal Tree and the other locally adaptive variants (LWCI, CQR), where Conformal Tree achieves the minimal average interval width on datasets Data 1 and star, and is competitive with the best method on almost all datasets. It is important to note that the other locally adaptive variants (LWCI and CQR) require an auxiliary model fit on the training data. While Conformal Tree may not always outperform CQR, unlike CQR it is applicable when the practitioner only has access to the calibration data.
@@ 5.1 Classifying States of Legislators using an LLM
We apply our Conformal Tree methodology, partitioning the two-dimensional space of DW-NOMINATE scores using a dyadic tree that captures local variations in the conformity score. Our conformal prediction sets have different probability thresholds in each tree leaf, causing local variations in the sizes of the conformal sets that reflect the variation in the calibration data. This type of variation reflects quality (informativeness) of prompts for uncertainty quantification. In our experiments, we set $m=200$ and $K=20$.
@@ 5.2 Conformalizing ChatGPT Diagnoses of Skin Diseases
Services hosting black-box generative language models like ChatGPT have now become widespread, offering powerful tools for applications ranging from customer support to creative content generation. As a result, these models are being increasingly integrated into everyday applications, making sophisticated AI tooling accessible to the general public. One significant yet controversial use lies in healthcare, where people have the option to use ChatGPT to obtain preliminary diagnoses of their medical conditions based on physical descriptions and perceived symptoms [11]. Previous studies have found ChatGPT to be effective at creating shortlists of possible diagnoses based on clinical vignettes from case reports [32] and to be an effective self-diagnostic tool for common orthopedic diseases [15]. Uncertainty quantification of these black-box generative models is important for its potential future use in medical diagnostics, but also addresses the current reality that people may use these tools for self-diagnosis, despite potential risks.
@@
Figure 5: Obtaining diagnoses by asking a large language model. Exact list of covariates and prompt used are available in Section C.
@@
We applied Conformal Tree to obtain 90% prediction sets for erythemato squamous disease diagnoses based on patient characteristics and physical descriptions. Erythemato squamous disease is a class of common skin diseases, that can manifest as rashes, redness, or scaly patches in the skin. The disease is primarily categorized in six different forms. We consider an example of using generative AI for self-diagnosis by applying GPT-4o to give diagnoses for the type of erythemato squamous disease present based only on physical characteristics that a patient could report from an inspection. The process by which self-diagnoses are obtained from the language model is outlined in Figure 5, with full details given in Section C. On a heldout set of 20% of the data, the disease which ChatGPT assigns highest probability is the correct disease 51% of the time, which is significantly better than a random classifier baseline, which guesses each class with probability proportional to its underlying probability in the data, which achieves only a 20% accuracy. However, as one might expect, ChatGPT is outperformed by standard machine learning tools, as a random forest model achieves 85% accuracy on the same heldout test data. We observe that some signal is present inside ChatGPT’s responses, but it is outperformed by a model specifically trained for this task, as should be expected.
In order to calibrate the data, we used a dataset of clinical evaluations of 12 features from patients, comprised of a mixture of continuous, ordinal, and binary variables. We used 80% of the data to calibrate our method, and heldout the remaining 20% in order to test it. Once the diagnoses yielded from the model are in hand, Conformal Tree provides a way to quantify the LLM’s uncertainty when classifying the disease, yielding a prediction set that will contain the true eventual biopsy result with 90% probability. For our experiment, we set $m=10$ and $K=20$.
Figure 6: (Left): Comparison of prediction set sizes for erythemato squamous disease using GPT-4o. Conformal Tree yields prediction sets that are smaller than standard conformal prediction, containing 1.33 less items on average. (Right) Class-wise average prediction set sizes for Conformal Tree and standard split-conformal prediction.
@@
Figure 7: First five levels of Conformal Tree’s structure on the skin disease calibration data. Node colors correspond to the mid-range of conformity scores within the node. Below each node, we plot the empirical distribution of calibration data corresponding to each disease within that node. See Figure 9 for an analogous plot visualizing the heldout test data.
The partition created by Conformal Tree found that GPT is relatively better at classifying the disease in cases where the patient reports high amounts of erythema (redness), scaling, and itching, the disease has definite borders, the koebner phenomenon is present, and the papules are not polygonal. This clinical combination of attributes seems to place large probability mass on psoriasis. On the other hand, when erythema and scaling and present, but the borders are indefinite and there is no scaling, GPT creates the largest conformal sets, as there is insufficient information for it to meaningfully distinguish between the different diseases. In Figure 6 (a), we compare the performance of Conformal Tree to standard split-conformal prediction in terms of the 90% prediction set sizes on the heldout test data. We observe that Conformal Tree results in 90% prediction sets that contain 1.33 less entries on average, and have significantly more test sets with only a single entry. We also note an average of 2.05 less entries compared to naive uncertainty quantification (Section 5.1.1). Over 93% of the 90% prediction sets from Conformal Tree are as large or smaller than those of standard split-conformal prediction. Figure 6 (b) shows the class-wise average prediction set sizes for Conformal Tree compared to standard split-conformal prediction. We observe that Conformal Tree yields much smaller prediction sets for lichen planus and chronic dermatitis. In Figure 8 in Section C.3, we show empirical test coverage for split-conformal and Conformal Tree for varying $1-\alpha$ values on the GPT diagnosis example. Conformal tree achieves the desired coverage level across $\alpha$ values, and does not overcover.

### v2

@@ 2.2 Robust Regression Trees
(Conformal Forests) Gasparin and Ramdas [17] have recently proposed a method for combining conformal sets via a weighted majority vote procedure which preserves a weakened coverage guarantee. We briefly describe how this can be used to extend Conformal Tree to the random forest setting. One could grow $h$ different trees, for example, by subsampling the calibration data, randomizing splitting dimensions, or both. The prediction set corresponding to each tree can be combined using the weighted majority vote procedure [17], yielding an aggregated interval. When applied to our robust dyadic trees in Algorithm 2, the aggregated interval will have marginal coverage with probability at least $1-2\alpha-2\delta$. Such an aggregated interval may sacrifice conditional coverage in each leaf node, as well as some of the coverage guarantee, in exchange for a smoother and more locally adaptive conformal set. As dyadic trees split always at the center, we could apply (random) rotation of the covariates when fitting multiple trees similar to [59] to diversify the tree fits. Such an approach will require a device of integrating the local weighting mechanism in [59] with the the weighted majority vote procedure [17]. In our experiments (Table 2), we show simply the random forest version of conformal tree, using the regular CART with the majority vote in [17].
@@ 3 Theoretical Underpinnings
Let $(X_{1},Y_{1}),\ldots,(X_{n+1},Y_{n+1})$ be i.i.d. samples from a joint probability measure $\mathbb{P}$. Let $\widehat{\mathfrak{X}}(\mathcal{D}_{1:j})$ denote the partition resulting from Algorithm 2 on data $(X_{1},S(X_{1},Y_{1})),\ldots,$ $(X_{j},S(X_{j},Y_{j})))$ for $j\in\{n,n+1\}$. Denote by $m$ the hyperparameter limiting the minimum number of samples in each leaf node and $K_{\rm max}<(n+1)/m$ be the maximum number of leaf nodes. Define
$\delta(n,m):=\frac{2}{m}+\exp\left\{-\left(\frac{n+1}{K_{\rm max}}-m\right)\right\}$ (3.1)
@@
Let $(X_{1},Y_{1}),\ldots,(X_{n+1},Y_{n+1})$ be i.i.d. samples from a joint probability measure $\mathbb{P}$. Denote by $\widehat{C}_{n,m}(X_{n+1})$ the prediction interval on $X_{n+1}$ constructed by Algorithm 1, with a minimum of $m$ samples in each leaf. Denote the obtained partition by $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})=\{\mathscr{X}_{k}\}_{k=1}^{K}.$ Then, we have for any $k\in\{1,\dots,K\},$
$\mathbb{P}\{Y_{n+1}\in\widehat{C}_{n,m}(X_{n+1})\mid X_{n+1}\in\mathscr{X}_{k},\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})\}\geq 1-\alpha-\delta(n,m)\,$
$\mathbb{P}\{Y_{n+1}\in\widehat{C}_{n,m}(X_{n+1})\mid X_{n+1}\in\mathscr{X}_{k},\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})\}\leq 1-\alpha+\frac{1}{1+m}+\delta(n,m)\,.$ (3.3)
@@
Let $(X_{1},Y_{1}),\ldots,(X_{n+1},Y_{n+1})$ be i.i.d. samples from a joint probability measure $\mathbb{P}$. Denote by $\widehat{C}_{n,m}(X_{n+1})$ the prediction interval on $X_{n+1}$ constructed by Algorithm 1, with a minimum of $m$ samples in each leaf. Then, we have
@@
$\mathbb{P}\{Y_{n+1}\in\widehat{C}_{n,m}(X_{n+1})\}\leq 1-\alpha+\frac{1}{1+m}+\delta(n,m)\,.$ (3.5)
@@
While the bounds in Theorem 3.2 are not entirely optimistic as we upper-bound the probability of miscoverage by 1 when $\widehat{\mathfrak{X}}(\mathcal{D}_{1:n})\neq\widehat{\mathfrak{X}}(\mathcal{D}_{1:n+1})$. We found empirically, however, that the Conformal Tree yields the desired coverage on most examples we tried (as summarized in Table LABEL:table:all-results). One may also restore the originally intended coverage level $1-\alpha$ by using the $1-\alpha+\delta(n,m)$ quantile (if below one) to define the conformal interval within each leaf region.
@@ 4 Simulation and Benchmark Studies
In this section, we consider regression tasks on simulated and real benchmark data. We use the absolute residuals as the conformity score. As our primary evaluation metric, we compare the average length of the prediction interval and proportion of test intervals smaller than that of split conformal (proportion better, or in short, P.B.). We also measure empirical coverage rates on the test dataset. In Section C, a sensitivity analysis for the hyperparameters $K_{\rm max}$ and $m$ is provided. Additional information on data sets, predictive models, and experiments can be found in Supplement D.
@@ 4.1 Conformal Tree is Adaptive
These two setups represent heteroskedasticity for distinct reasons. Data 1 has the noise standard deviation scaling with $X$, causing the distribution of residuals to vary across the domain. Therefore, any single regression (mean) model cannot avoid increasing errors for the larger covariate due to the increasing variation of the data. Data 2 is difficult to model near $X=0$ due to the $X^{-3}$ term inside the $\sin$, causing rapid change in the mean function that requires a lot of data to model accurately. Therefore, due to the lack of model fit, the error is large for the smaller covariate values.
Figure 1: Conformal Tree can adaptively reflect locality without an auxiliary model of quantiles or conformity scores fit on the training dataset. We display the average width, empirical coverage on held-out test data, and proportion better. The interval corresponds to a single random trial among the 10 random trials.
We train a regression model $\hat{\mu}$ a fully-connected deep feed forward neural network with a smooth regularization penalty. The trained model behaves as the orange line in Figure 1. We synthetically generated datasets of size 1000 and randomly split the data into equally sized calibration and test datasets ten times, and applied each method for each split. We apply our Conformal Tree method ($m=100$ and maximum leaves as $5$) and compare it with the standard split-conformal prediction. Figure 1 highlights that Conformal Tree can adaptively reflect local changes without the necessity for an auxiliary model of quantiles or conformity scores on the training dataset. As the partitioning method is robust to the addition of a new iid data point, the obtained prediction intervals with $\alpha=0.1$ still have good coverage at 0.899 for Data 1 and 0.904 for Data 2 when averaged over 10 random trials. For the standard conformal prediction, the average coverage was 0.904 and 0.874, respectively. Furthermore, the average percentage of points for which the interval was narrower is 0.54 for Data 1 and 0.5 for Data 2.
@@ 4.2 Benchmark Datasets
We consider various real data examples that had previously been used in the literature for comparing the efficacy of regression prediction intervals [44, 47, 42], namely, bike, bio, community, concrete, homes, and star. The number of observations in these datasets vary from 200 to about 50,000, and the number of predictor variables ranges from 1 to 100. We also consider the simulation data (Data 1 and Data 2) in the previous section. We randomly split each dataset into training, test, and calibration sets, with proportions 0.3, 0.2, and 0.5 respectively. On the training dataset, a random forest model is fitted as the “black-box” predictive model $\hat{\mu}$. For each dataset, we repeat the entire procedure 5 times, including splitting the data and forming the intervals. We report the average metrics taken over these trials.
We compare with standard split-conformal prediction [26], locally-weighted conformal inference (Locally Weighted) [26], and conformalized quantile regression (CQR) [42]. Locally-weighted conformal inference is similar to our method in that it relies on a locally varying model for the conformity score $\hat{\sigma}:\mathcal{X}\to\mathbb{R}$, which is fit on the training data. We use CART as well as a random forest model for $\hat{\sigma}$ in our experiments. When applying CQR, we likewise employ two versions, where the model for the conditional upper and lower quantiles is a single tree and a forest with 100 trees respectively. Recall that our setting considers the inaccessibility of the original training data used for training the regression model. Since all of these benchmark algorithms need access to the training data set (or some other set other than the calibration set) to train their auxiliary or grouping functions, we used a random half of the calibration split to train the auxiliary functions and the remaining half for calibrate.
Split Conformal Conformal Tree Locally Weighted (Tree) CQR (Tree)
Width Cov. ISL Width Cov. ISL P.B. Opt. $m$ Opt. $K$ Width Cov. ISL P.B. Width Cov. ISL P.B.
Data 1 ($n=499$) 4.29 0.92 5.28 3.5 0.86 4.66 0.59 20 4 3.52 0.88 4.48 0.54 4.84 0.93 5.77 0.0
Data 2 ($n=499$) 2.93 0.91 6.27 2.96 0.86 3.98 0.89 20 8 3.53 0.91 4.39 0.54 8.4 0.94 10.23 0.0
bike ($n=5443$) 1.37 0.9 2.17 1.33 0.88 2.06 0.61 20 64 1.22 0.89 1.52 0.67 2.67 0.89 4.15 0.0
bio ($n=22864$) 1.7 0.9 2.27 1.66 0.89 2.24 0.67 20 64 1.63 0.9 2.25 0.57 3.37 0.89 3.93 0.0
community ($n=997$) 1.9 0.89 2.89 1.78 0.89 2.8 0.55 20 32 1.77 0.9 2.46 0.59 3.6 0.9 5.09 0.0
concrete ($n=514$) 0.68 0.92 0.85 0.61 0.89 0.79 0.72 20 8 0.7 0.93 0.84 0.21 1.47 0.91 1.78 0.0
homes ($n=10807$) 0.62 0.91 1.18 0.64 0.9 1.03 0.7 20 32 0.7 0.9 0.95 0.63 1.19 0.9 2.18 0.0
star ($n=1080$) 0.17 0.89 0.22 0.17 0.87 0.23 0.67 20 4 0.17 0.9 0.23 0.21 0.29 0.91 0.33 0.0
Table 1: Empirical results on real and synthetic data for tree-based localization methods for $\alpha=0.1$. We display the average over five trials of the width, coverage, interval score length (ISL) and proportion of test points with intervals smaller than split conformal (P.B.). We display the calibration set size as $n$. Hyperparameters $m$ and ${K_{\rm max}}$ are tuned for each dataset and for each method by minimizing ISL on held-out validation data over a coarse grid.
In addition, for all methods, we further partition one fifth of the remaining calibration set for each method to be used to tune hyperparameters. We tune the minimum samples per leaf, the maximum leaf nodes, and the number of trees (when applicable) for each method, by selecting the combination of hyperparameters that yield the minimal interval score loss [19] when evaluated on the held-out tuning data. Interval score loss (ISL) with threshold $\alpha$ is a proper scoring rule for prediction intervals that is minimized by the true $\alpha/2$ and $21-\alpha/$ 2 conditional quantiles [19] and has been used to evaluate the performance of conformal prediction methods [44]. The ISL for observation $y$ given the $1-\alpha$ prediction interval $[\ell,u]$ is given by $(u-\ell)+\frac{2}{\alpha}(\ell-y)\mathbb{I}\{y<\ell\}+\frac{2}{\alpha}(y-u)\mathbb{I}\{y>u\}$. In Appendix E, we also compare with other similar methods involving the use of regression trees and forests such as Izbicki et al. [23], Cabezas et al. [6], Martinez et al. [33], which find data-driven groups on additional datasets as well.
Split Conformal Conformal Forest Locally Weighted (Forest) CQR (Forest)
Width Cov. ISL Width Cov. ISL P.B. Width Cov. ISL P.B. Width Cov. ISL P.B.
Data 1 ($n=499$) 4.2 0.89 5.8 4.12 0.95 4.57 0.51 3.71 0.92 4.53 0.48 3.67 0.95 4.18 0.65
Data 2 ($n=499$) 2.81 0.87 6.22 3.05 0.86 4.09 0.91 3.1 0.86 5.31 0.71 4.06 0.86 5.38 0.76
bike ($n=5443$) 1.38 0.9 2.15 1.14 0.93 1.28 0.72 1.1 0.9 1.64 0.7 1.46 0.89 1.62 0.49
bio ($n=22864$) 1.71 0.9 2.27 1.52 0.92 1.95 0.62 1.43 0.9 2.02 0.69 1.74 0.89 1.98 0.33
community ($n=997$) 1.84 0.88 2.97 1.64 0.9 2.39 0.66 1.66 0.91 2.41 0.63 1.67 0.9 2.39 0.59
concrete ($n=514$) 0.72 0.93 0.85 0.7 0.92 0.84 0.66 0.66 0.93 0.81 0.78 1.13 0.9 1.25 0.0
homes ($n=10807$) 0.63 0.91 1.25 0.64 0.93 0.85 0.69 0.57 0.9 0.89 0.74 0.76 0.9 1.01 0.6
star ($n=1080$) 0.18 0.92 0.2 0.19 0.94 0.2 0.23 0.18 0.92 0.2 0.46 0.18 0.92 0.21 0.36
Table 2: Empirical results on real and synthetic data for forest-based localization methods for $\alpha=0.1$. We display the average over five trials of the width, coverage, interval score length (ISL) and proportion of test points with intervals smaller than split conformal (P.B.). We display the calibration set size as $n$. Hyperparameters $m$ and ${K_{\rm max}}$ are tuned for each dataset and for each method by minimizing ISL on held-out validation data. We use 100 trees for all forest-based methods.
The results for $\alpha=0.1$ are displayed in Table 1, which clearly displays the tightening effect of Conformal Tree. It shows that the average interval widths and interval score loss values are competitive with the other tree-based methods. Moreover, Conformal Tree has the highest proportion of test points with intervals smaller than split conformal on all datasets except two (bike and community). Tables 1 also shows the average empirical coverage on the held-out test data for each method. We observe that each method nearly achieves the desired coverage level of 0.9. While our method shows slight undercoverage relative to the nominal level $1-\alpha$, it is still within our theoretical bounds of $1-\alpha-\delta$. In Table 2, we also display the results of a variant of our method that uses a random forest of 100 CART trees (with bagging and random feature selection) for partitioning. We aggregate these prediction sets using a majority vote procedure (See Remark 3). We find this method to yield the lowest interval score length among forest-based procedures on the majority of datasets in our comparison, excluding Data 1 and concrete. Recall that this procedure only promises coverage up to $1-2\alpha-2\delta$, though we see coverage above 0.9 in all but one of our datasets. We note that this method sacrifices interpretability in exchange for a more flexible model for the conditional conformity score.
@@ 5.1 Classifying States of Legislators using an LLM
We apply our Conformal Tree methodology, partitioning the two-dimensional space of DW-NOMINATE scores using a dyadic tree that captures local variations in the conformity score. Our conformal prediction sets have different probability thresholds in each tree leaf, causing local variations in the sizes of the conformal sets that reflect the variation in the calibration data. This type of variation reflects quality (informativeness) of prompts for uncertainty quantification. In our experiments, we set $m=200$ and ${K_{\rm max}}=7$, where seven is the largest value possible within the restriction ($K_{\rm max}<(n+1)/m$).
@@ 5.2 Conformalizing ChatGPT Diagnoses of Skin Diseases
One significant yet controversial use of ChatGPT lies in healthcare, where people have the option to use it to obtain preliminary diagnoses of their medical conditions based on physical descriptions and perceived symptoms [16]. Previous studies have found ChatGPT to be effective at creating shortlists of possible diagnoses based on clinical vignettes from case reports [48] and to be an effective self-diagnostic tool for common orthopedic diseases [25]. Uncertainty quantification of these black-box generative models is important for its potential future use in medical diagnostics, but also addresses the current reality that people may use these tools for self-diagnosis, despite potential risks.
@@
Figure 5: Obtaining diagnoses by asking a large language model. Exact list of covariates and prompt used are available in Section D.2.1.
@@
We consider the erythemato squamous disease diagnosis task based on patient characteristics and physical descriptions [22], which consists of 366 observations, 35 features. Erythemato squamous disease is a class of common skin diseases, that can manifest as rashes, redness, or scaly patches in the skin. The disease is primarily categorized in six different forms. We consider an example of using generative AI for self-diagnosis by applying GPT-4o to give diagnoses for the type of erythemato squamous disease present based only on physical characteristics that a patient could report from an inspection. The process by which self-diagnoses are obtained from the language model is outlined in Figure 5, with full details given in Section D.2.1. We used the 12 features from patients, comprised of a mixture of continuous, ordinal, and binary variables. We ignore the remaining 23 histopathological features, as the purpose of this experiment is to emulate a user describing physical symptoms to a language model.
Figure 6: (a): Comparison of prediction set sizes for erythemato squamous disease using GPT-4o. (b) Class-wise prediction accuracy of ChatGPT. (c) The prediction confidence of ChatGPT on the true label. The ChatGPT performance order among labels in (b) and (c) aligns with the tightness order in (a) of Conformal Tree results.
On a heldout set of 20% of the data, the disease which ChatGPT assigns highest probability is the correct disease 51% of the time, which is significantly better than a random classifier baseline, which guesses each class with probability proportional to its underlying probability in the data, which would achieve only a 1/6 accuracy. However, as one might expect, ChatGPT is outperformed by a model specifically trained for this task, as a random forest model achieves 85% accuracy on the same heldout test data. We applied Conformal Tree to obtain 90% prediction sets with $m=10$ and ${K_{\rm max}}=20$, as well as the Split Conformal prediction and Naive UQ. For conformal predictions, we used 60% of the data to calibrate our method, and heldout the remaining 40% in order to test it. As the prediction probability vectors from ChatGPT tend to be limited to a certain set of values, we added a tiny random perturbation of uniform(0.0001, 0.001) to break the tie and normalized (see, Tibshirani et al. [53]).
Figure 6 (a) shows the class-wise average prediction set sizes for Conformal Tree compared to standard split-conformal prediction. To explain the interpretability of this result, we also show additional information in Figures 6 (b) and (c), which represent the performance of ChatGPT in terms of prediction accuracy and confidence, respectively. Now, we can see that the predicted set size in Figure 6 (a) is smallest for psoriasis and then for lichen planus, seboreic dermatitis, and chronic dermatitis, which aligns the order of the prediction accuracy and confidence observed in Figure 6 (b) and (c). In terms of overall prediction efficiency, we draw the frequency plot in Figure 9 in Section E, which shows the frequency of a single element set is the highest for Conformal Tree.
@@
Figure 7: Conformal Tree’s full structure on the skin disease conformal data, with $m=50$ and $K_{max}=4$. Node colors correspond to the mid-range of conformity scores (of calibration data) within the node. Below each node, we plot the empirical distribution of calibration data corresponding to each disease within that node. The tree and partition correspond to that of Figure 10 in Section E.
For a more interpretable presentation of the results, we intentionally set $m=50$ to obtain a tree with a smaller number of partition boxes. In Figure 7, we visualize each level of Conformal Tree’s structure on the skin disease calibration data. This plot emphasize the interpretability of our method by showing the precise splits that determine each leaf node, in which group conditional coverage is achieved. In Figure 10 in Section E, we visualize the empirical coverage and threshold for decreasing $\alpha$ on each of the resulting partitions, obtained by applying Conformal Tree on one realization of the split of the data. The result is very interpretable. In Box 3, the threshold values of our method (orange color) are significantly smaller than those of split conformal prediction. Interestingly, Box 3 does not contain any data points of pityriasis rosea and pityriasis rubra pilaris, for which the ChatGPT shows the lowest performance. On the other hand, in Box 1, the thresholds of our method are set large even for the largest $\alpha$ (and increasing slightly as $\alpha$ decreases), as the pityriasis groups are heavily present in this box. The estimated coverage in Figure 10 can be improved if we had more test data points. It is implied by the dramatic difference in Figure 11 (a) and (b) in Section E that coverage estimation requires a large number of test data points. In Figure 11, we show empirical test coverage for split-conformal and Conformal Tree for varying $1-\alpha$ values in the GPT diagnosis example. When the size of data is artificially increased by multiple random splits (see Section E for details), the estimation quality improves. Note that, however, even without multiple splits (Figure 11 (a)), the estimated coverage for conformal tree achieves the desired coverage level across the values of $\alpha$, staying within the theoretical bounds. Note that our methodology does not guarantee class-wise coverage for each true class but box-wise coverage. For the tree structure of $m=20$, we also provide visualization of the interpretable tree-splitting process in Figure 12 in Section E.